OpenStudy (seattle12345):

can some one please help with this. Sqrt(x+1)=x-1

OpenStudy (anonymous):

\[\sqrt{x} ^{2} =x\]

OpenStudy (anonymous):

is it like this?  if so square both sides and get x+1 = (x-1)^2 (x-1)^2 = x^2 - 2x + 1 so x^2 - 2x + 1 = x + 1 x^2 - 3x = 0 factor out x x(x-3) = 0 so x = 0 and x = 3

OpenStudy (seattle12345):

@jay657 yes.

OpenStudy (anonymous):

yea so your answer should be x = 0 and x = 3 as i stated above

OpenStudy (anonymous):

\[\sqrt{x+1} = x-1\] Squaring both sides: \[x+1 = (x-1)^2\] \[x+1 = x^2 + 1 - 2x\] \[x^2 - 3x = 0\] \[x(x-3) = 0\] So, \(x = 0\) And \(x = 3\)..

OpenStudy (anonymous):

Any problem @seattle12345 ??

OpenStudy (seattle12345):

@waterineyes I guess I didnt realize it needed to be factored. Is that because it is quadratic?